# Model-independent predictions for decays of double-heavy hadrons into pairs of heavy hadrons

### v2

@@ III.2.2 Selection Rules for $\Sigma_{\eta}^{\epsilon}$
If the two static hadrons are identical (for $(\bm{3},\bm{3})$ sources) or charge conjugates (for $(\bm{3},\bm{3}^{\ast})$ sources), the selection rule for $\epsilon$ in Eq. (47) implies a selection rule for $\eta$. Conservation of $\eta$, Eq. (8), and conservation of $\epsilon$, Eq. (47) with $\pi_{2}=\pi_{1}$ or $\pi_{1}(-1)^{2j_{1}}$, together imply
@@
Concretely, Eq. (48) states that a light-QCD state with B-O quantum numbers $\Sigma_{g}^{-}$ or $\Sigma_{u}^{+}$ in the presence of $(\bm{3},\bm{3})$ sources cannot decay into a pair of identical static hadrons. Similarly, a light-QCD state with B-O quantum numbers $\Sigma_{g}^{-}$ or $\Sigma_{u}^{+}$ in the presence of $(\bm{3},\bm{3}^{\ast})$ sources cannot decay into a pair of charge-conjugate static hadrons.
@@ III.2.3 Selection Rules for Orbital Angular Momentum and Spin
There are selection rules associated with conservation of parity and angular momentum. For a double-heavy hadron with parity $P$, conservation of parity implies
@@
III.2.4 Selection Rules from Identical Particles
For two identical heavy quarks, there is a selection rule due to fermionic signs that constrains the values of $L_{Q}$ and $S_{Q}$. Exchanging their positions produces a sign $\eta(-1)^{L_{Q}}$, while exchanging their spin states produces a sign $(-1)^{S_{Q}+1}$. Exchanging the two identical heavy quarks must give an overall sign $-1$, because they are fermions. The constraint from these signs is
$(-1)^{L_{Q}+S_{Q}}=\eta.$ (52)
For two identical heavy hadrons, which requires $J_{2}=J_{1}$, there is a selection rule that constrains the values of $L_{Q}^{\prime}$ and $S$. Exchanging their positions produces a sign $(-1)^{L_{Q}^{\prime}}$, while exchanging their spin states produces a sign $(-1)^{S+2J_{1}}$. Exchanging the two identical heavy hadrons must give an overall sign $(-1)^{2J_{1}}$, which is equal to $+1$ if they are bosons with integer $J_{1}$ and $-1$ if they are fermions with half-integer $J_{1}$. The constraint from these signs is
$(-1)^{L_{Q}^{\prime}+S}=+1.$ (53)
@@ III.3 Relative Partial Decay Rates
The coupling potentials defined by inserting Eq. (41) into Eq. (43) depend on the $QQ$ or $Q\bar{Q}$ distance $r$ only through the B-O transition amplitudes $g_{\lambda,\eta}$ introduced in Eq. (25). In general, calculating the decay width of a double-heavy hadron into a pair of heavy hadrons requires the solution of the coupled-channel Schrödinger equation. However, in some simple cases one can derive analytic expressions for the relative partial decay rates without solving any Schrödinger equation.
Let us consider the decays of two double-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ into heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$, respectively. The double-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ belong to the same B-O multiplet with angular-momentum/parity $L^{P}$ and they have total heavy spins $S_{Q\alpha}$ and $S_{Q\beta}$ and total hadron spins $J_{\alpha}$ and $J_{\beta}$. The heavy hadrons $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}}$ belong to the same HQSS doublet with light-QCD quantum numbers $j_{1}^{\pi_{1}}$ and they have spins $J_{1\alpha}$ and $J_{1\beta}$. The heavy hadrons $\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ belong to the same HQSS doublet with light-QCD quantum numbers $j_{2}^{\pi_{2}}$ and they have spins $J_{2\alpha}$ and $J_{2\beta}$.
We consider decays into heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ with the same orbital angular momentum $L_{Q}^{\prime}$. Their coupling potentials with the double-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ are obtained by inserting Eq. (41) into Eq. (43). Note that, in addition to the orbital angular momentum $L_{Q}^{\prime}$, the coupling potentials also depend on the static angular momentum $S$, which for the heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ or $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ is just the total heavy-hadron spin $S_{\alpha}$ or $S_{\beta}$. We will assume here that the angular distributions of the heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ are not measured. We therefore consider decay rates summed over all possible values of $S_{\alpha}$ and $S_{\beta}$.
@@
The simple ratio of coupling potentials implies a simple ratio of partial decay rates if in addition the kinetic energies of the heavy hadrons $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}}$, $\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}}$, $\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ are much larger than their spin splittings, that is, if the masses of the double-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ are both well above the corresponding hadron-pair thresholds. In this case, the ratio of the partial decay rates normalized by the phase-space factors $v_{\alpha}^{2L_{Q}^{\prime}+1}$ and $v_{\beta}^{2L_{Q}^{\prime}+1}$ can be approximated by the square of the ratio of the coupling potentials, which is a rational number:
@@
where $N_{\alpha}$ and $N_{\beta}$ are normalization coefficients defined after Eq. (43). The factor $N_{\alpha}^{2}$ is 1 unless the two static hadrons $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$ are identical (for $(\bm{3},\bm{3})$ sources) or charge conjugates (for $(\bm{3},\bm{3}^{\ast})$ sources). Even in this case, $N_{\alpha}^{2}$ differs from 1 only if $J_{1\alpha}\neq J_{2\alpha}$, in which case $N_{\alpha}^{2}=2$. If the two double-heavy hadrons are the same, which requires $S_{Q\alpha}=S_{Q\beta}$ and $J_{\alpha}=J_{\beta}$, Eq. (54) gives a branching ratio for the double-heavy hadron $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$.
Note that the relative partial decay rates in Eq. (54) are independent of the orbital angular momentum $L_{Q}$ of the $QQ$ or $Q\bar{Q}$ pair inside the double-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$. The dependence of the coupling potential $V^{J,P,\eta^{\prime}}_{S_{Q},L,\eta}$ in Eq. (43) on $L_{Q}$ enters only through the $G_{L,\eta}^{P}$ factor defined in Eq. (41). If only a single $g_{\lambda,\eta}$ factor contributes to the decay, then only a single $G_{L,\eta}^{P}$ contributes and it cancels in the relative partial decay rates. Not depending on $L_{Q}$ is an essential condition for the relative partial decay rates to be predictable without knowing the radial dependence of the B-O potentials. In fact, a double-heavy-hadron state may contain several components with different values of $L_{Q}$ since it is in general not conserved. The relative partial decay rates in Eq. (54) do depend on the B-O angular-momentum $L$, which is conserved in the heavy-quark limit.
@@ IV Quarkonia and Quarkonium Hybrids
We will only consider decays into the three lightest pairs of heavy mesons: $S+S$, $S+P^{-}$, and $S+P^{+}$. For the sake of simplicity, we choose the first heavy meson ($S$-wave) to have the flavor of a light antiquark and the second heavy meson ($S$-, $P^{-}$-, or $P^{+}$-wave) to have the flavor of a light quark. The $\mathcal{C}\mathcal{P}$ eigenvalue $\eta^{\prime}$ of the meson pair is determined by the total heavy spin $S_{Q}$ and B-O quantum number $\eta$ of the double-heavy hadron as $\eta^{\prime}=(-1)^{S_{Q}+1}\eta$ as given in Eq. (44).
@@ IV.1 Quarkonium Decays
Quarkonia are associated with bound states in the lowest confining $\Sigma_{g}^{+}$ potential, with B-O quantum numbers $\lambda=0$, $\eta=+1$, and $\epsilon=+1$. Their light-QCD state at short distances is dominated by angular-momentum/parity $j^{\pi}=0^{+}$. The B-O angular momentum is therefore equal to the orbital angular momentum: $L=L_{Q}$. A quarkonium multiplet consists of several quarkonia with same orbital angular momentum $L_{Q}$ and parity $P=(-1)^{L_{Q}+1}$ but different quantum numbers $S_{Q}$ and $J$. The $C$-parity is determined by the orbital angular momentum $L_{Q}$ and the total heavy spin $S_{Q}$ as $C=(-1)^{L_{Q}+S_{Q}}$.
@@ IV.2 Quarkonium Hybrid Decays
The lowest-energy quarkonium hybrids are associated with bound states in confining $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials. Their light-QCD state at short distances is dominated by angular-momentum/parity $j^{\pi}=1^{+}$. A hybrid multiplet consists of several quarkonium hybrids with the same B-O angular momentum $L$ and parity $P$ but different quantum numbers $S_{Q}$ and $J$. The $C$-parity is determined by the parity $P$ and the total heavy spin $S_{Q}$ as $C=P(-1)^{S_{Q}}$. There are three different types of hybrid multiplets:
@@ V.1 Transition Amplitudes
where the $\xi$’s are angular-momentum coefficients that are the same for all nonflip triplet decay models and the $A$’s are matrix elements involving the spatial overlap of quark-model wavefunctions that depend on the specific model. Their dependence on the quantum numbers associated with the initial and final mesons is denoted by arrays of numbers in square brackets. The sums in Eq. (63) are over quantum numbers $L^{\prime}$ and $L^{\prime\prime}$ associated with the orbital-angular-momentum vectors $\bm{L}^{\prime}\equiv\bm{L}_{1}+\bm{L}_{2}$ and $\bm{L}^{\prime\prime}\equiv\bm{L}^{\prime}+\bm{L}_{Q}^{\prime}$. The subscripts $\pm$ in Eq. (63) stand for two different “topologies” of the decay. The interference between the (+)(+)( + ) and ()(-)( - ) topologies ensures conservation of $C$-parity. The angular-momentum coefficients can be written explicitly in terms of Wigner 6-$j$ and 9-$j$ symbols as [38]
@@ V.2 Selection Rules
The transition amplitude for the decays of a meson into pairs of mesons are obtained by inserting Eq. (64) into Eq. (63). Zeroes of the transition amplitude correspond to selection rules. Zeroes of the angular-momentum coefficients $\xi$ correspond to selection rules associated with conservation of angular momentum. Zeroes of the spatial matrix elements $A$ correspond to selection rules associated with conservation of parity. These selection rules are described in Ref. [23].
For decays into meson pairs with identical spatial wavefunctions, there is an additional selection rule that corresponds to a cancellation in the sum of the two topologies (+)(+)( + ) and ()(-)( - ) in Eq. (63). It was derived in its most general form in Ref. [34] using the symmetries of the matrix elements $A$ for nonrelativistic quark models. The rule states that a quark-antiquark meson is forbidden from decaying into pairs of mesons with identical spatial wavefunctions if the $CP$-parity of the decaying meson and its total constituent-quark spin $S_{Q}$ stand in the relation $CP=(-1)^{S_{Q}+S_{q}+1}$, where $S_{q}$ is the total spin of the light-quark pair that is created. Since for nonflip triplet decay models one has $S_{q}=1$, the selection rule becomes
$CP=(-1)^{S_{Q}}.$ (65)
According to this selection rule, decays into pairs of mesons with identical spatial wavefunctions are allowed for quarkonia, which have $CP=(-1)^{S_{Q}+1}$, and forbidden for quarkonium hybrids, which have $CP=(-1)^{S_{Q}}$.
Quarkonium hybrids in B-O multiplets associated with bound states in the $\Pi_{u}$ potential, the $\Sigma_{u}^{-}$ potential, or coupled $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials, have $CP=(-1)^{S_{Q}}$. The B-O equivalent of two quark-model mesons with identical spatial wavefunctions is two heavy mesons in charge-conjugate HQSS doublets, e.g., two $S$-wave mesons, two $P^{-}$-wave mesons, or two $P^{+}$-wave mesons. Our finding that quarkonium hybrids in B-O multiplets with a $\Sigma_{u}^{-}$ component may have nonzero decay rates into pairs of $S$-wave mesons therefore contradicts the selection rule in Eq. (65), as first pointed out in Ref. [27]. This observation can be generalized to meson pairs other than $S+S$ by applying the selection rules in Section III.2. Bound states in the $\Sigma_{u}^{-}$ potential or coupled $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials can have nonzero decay rates into pairs of heavy mesons in charge-conjugate HQSS doublets with $j_{1}=j_{2}\geq\frac{1}{2}$, such as $S+S$, $P^{-}+P^{-}$, or $P^{+}+P^{+}$. Bound states in the $\Pi_{u}$ potential can have nonzero decay rates into pairs of heavy mesons in charge-conjugate HQSS doublets with $j_{1}=j_{2}\geq\frac{3}{2}$, such as $P^{+}+P^{+}$.

### v3

@@ III.2.2 Selection Rules for $\epsilon$
If the two static hadrons are charge conjugates, the selection rule for $\epsilon$ in Eq. (36) implies a selection rule for $\eta$. Conservation of $\eta$, Eq. (8), and conservation of $\epsilon$, Eq. (36) with $\pi_{2}=\pi_{1}(-1)^{2j_{1}}$, together imply
@@
Concretely, Eq. (37) states that a hidden-heavy hadron with B-O quantum numbers $\Sigma_{g}^{-}$ or $\Sigma_{u}^{+}$ cannot decay into a pair of charge-conjugate static hadrons.
@@ III.2.3 Selection Rules for Orbital Angular Momentum and Spin
There are selection rules associated with conservation of parity and angular momentum. For a hidden-heavy hadron with parity $P$, conservation of parity implies
@@ III.3 Relative Partial Decay Rates
The coupling potentials defined by inserting Eq. (30) into Eq. (32) depend on the $Q\bar{Q}$ distance $r$ only through the B-O transition amplitudes $g_{\lambda,\eta}$ introduced in Eq. (25). In general, calculating the decay width of a hidden-heavy hadron into a pair of heavy hadrons requires the solution of a coupled-channel Schrödinger equation. However, in some simple cases one can derive analytic expressions for the relative partial decay rates without solving any Schrödinger equation.
Let us consider the decays of two hidden-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ into heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$, respectively. The hidden-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ belong to the same B-O multiplet with angular-momentum/parity $L^{P}$ and they have total heavy spins $S_{Q\alpha}$ and $S_{Q\beta}$ and total hadron spins $J_{\alpha}$ and $J_{\beta}$. The heavy hadrons $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}}$ belong to the same HQSS doublet with light-QCD quantum numbers $j_{1}^{\pi_{1}}$ and they have spins $J_{1\alpha}$ and $J_{1\beta}$. The heavy hadrons $\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ belong to the same HQSS doublet with light-QCD quantum numbers $j_{2}^{\pi_{2}}$ and they have spins $J_{2\alpha}$ and $J_{2\beta}$.
We consider decays into heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ with the same orbital angular momentum $L_{Q}^{\prime}$. Their coupling potentials with the hidden-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ are obtained by inserting Eq. (30) into Eq. (32). Note that, in addition to the orbital angular momentum $L_{Q}^{\prime}$, the coupling potentials also depend on the static angular momentum $S$, which for the heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ or $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ is just the total heavy-hadron spin $S_{\alpha}$ or $S_{\beta}$. We will assume here that the angular distributions of the heavy-hadron pairs $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}},\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}},\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ are not measured. We therefore consider decay rates summed over all possible values of $S_{\alpha}$ and $S_{\beta}$.
@@
The simple ratio of coupling potentials implies a simple ratio of partial decay rates if in addition the kinetic energies of the heavy hadrons $\Phi_{j_{1}^{\pi_{1}},J_{1\alpha}}$, $\Phi_{j_{2}^{\pi_{2}},J_{2\alpha}}$ and $\Phi_{j_{1}^{\pi_{1}},J_{1\beta}}$, $\Phi_{j_{2}^{\pi_{2}},J_{2\beta}}$ are much larger than their spin splittings, that is, if the masses of the hidden-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$ are both well above the corresponding hadron-pair thresholds. In this case, the ratio of the partial decay rates normalized by the phase-space factors $v_{\alpha}^{2L_{Q}^{\prime}+1}$ and $v_{\beta}^{2L_{Q}^{\prime}+1}$ can be approximated by the square of the ratio of the coupling potentials, which is a rational number:
@@
where $N_{\alpha}$ and $N_{\beta}$ are normalization coefficients defined after Eq. (32). The factor $N_{\alpha}^{2}$ is 1 unless the two static hadrons $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$ are charge conjugates. Even in this case, $N_{\alpha}^{2}$ differs from 1 only if $J_{1\alpha}\neq J_{2\alpha}$, in which case $N_{\alpha}^{2}=2$. If the two hidden-heavy hadrons are both $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$, which requires $S_{Q\beta}=S_{Q\alpha}$ and $J_{\beta}=J_{\alpha}$, the ratio in Eq. (41) gives a branching ratio for the hidden-heavy hadron.
Note that the relative partial decay rates in Eq. (41) are independent of the orbital angular momentum $L_{Q}$ of the $Q\bar{Q}$ pair inside the hidden-heavy hadrons $\Psi_{(L^{P},S_{Q\alpha})J_{\alpha}}$ and $\Psi_{(L^{P},S_{Q\beta})J_{\beta}}$. The dependence of the coupling potential $V^{J,P,\eta^{\prime}}_{S_{Q},L,\eta}$ in Eq. (32) on $L_{Q}$ enters only through the $G_{L,\eta}^{P}$ factor defined in Eq. (30). If only a single $g_{\lambda,\eta}$ factor contributes to the decay, then only a single $G_{L,\eta}^{P}$ contributes and it cancels in the relative partial decay rates. Not depending on $L_{Q}$ is an essential condition for the relative partial decay rates to be predictable without knowing the radial dependence of the B-O potentials. In fact, a hidden-heavy-hadron state may contain several components with different values of $L_{Q}$ since it is in general not conserved. The relative partial decay rates in Eq. (41) do depend on the B-O angular-momentum $L$, which is conserved in the heavy-quark limit.
@@
The relative partial decay rates in Eq. (41) apply to decays in which there is one single B-O transition amplitude $g_{\lambda,\eta}$. In instances where there is a dominant B-O transition amplitude $g_{\lambda,\eta}$ but there are also subdominant B-O transition amplitudes, the relative partial decay rates in Eq. (41) correspond to the leading-order approximation in which the subdominant B-O transition amplitudes are taken to zero. Ideally, small corrections to this approximation could be implemented using perturbation theory. But these corrections would spoil the factorization of the radial dependence that is necessary to predict the relative partial decay rates without solving a Schrödinger equation, making such a perturbative expansion of little practical utility.
@@ IV Quarkonia and Quarkonium Hybrids
We will only consider decays into the three lightest pairs of heavy mesons: $S+S$, $S+P^{-}$, and $S+P^{+}$. For the sake of simplicity, we choose the first heavy meson ($S$-wave) to have the flavor of a light antiquark and the second heavy meson ($S$-, $P^{-}$-, or $P^{+}$-wave) to have the flavor of a light quark. The $\mathcal{C}\mathcal{P}$ eigenvalue $\eta^{\prime}$ of the meson pair is determined by Eq. (33) in terms of the B-O quantum number $\eta$ of the hidden-heavy hadron and the total heavy spin $S_{Q}$.
@@ IV.1 Quarkonium Decays
Quarkonia are associated with bound states in the lowest confining $\Sigma_{g}^{+}$ potential, with B-O quantum numbers $\lambda=0$, $\eta=+1$, and $\epsilon=+1$. Their light-QCD state at short distances is dominated by angular-momentum/parity $j^{\pi}=0^{+}$. The B-O angular momentum is therefore equal to the orbital angular momentum, $L=L_{Q}$. A quarkonium multiplet consists of several quarkonia with same orbital angular momentum $L_{Q}$ and parity $P=(-1)^{L_{Q}+1}$ but different quantum numbers $S_{Q}$ and $J$. The $\mathcal{C}$-parity is determined by the orbital angular momentum $L_{Q}$ and the total heavy spin $S_{Q}$ as $C=(-1)^{L_{Q}+S_{Q}}$.
@@ IV.2 Quarkonium Hybrid Decays
The lowest-energy quarkonium hybrids are associated with bound states in confining $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials. Their light-QCD state at short distances is dominated by angular-momentum/parity $j^{\pi}=1^{+}$. A hybrid multiplet consists of several quarkonium hybrids with the same B-O angular momentum $L$ and parity $P$ but different quantum numbers $S_{Q}$ and $J$. The $\mathcal{C}$-parity is determined by the parity $P$ and the total heavy spin $S_{Q}$ as $C=P(-1)^{S_{Q}}$. There are three different types of hybrid multiplets:
@@ V.1 Transition Amplitudes
where the $\xi$’s are angular-momentum coefficients that are the same for all nonflip triplet decay models and the $A$’s are matrix elements involving the spatial overlap of quark-model wavefunctions that depend on the specific model. Their dependence on the quantum numbers associated with the initial and final mesons is denoted by arrays of numbers in square brackets. The sums in Eq. (50) are over quantum numbers $L^{\prime}$ and $L^{\prime\prime}$ associated with the orbital-angular-momentum vectors $\bm{L}^{\prime}\equiv\bm{L}_{1}+\bm{L}_{2}$ and $\bm{L}^{\prime\prime}\equiv\bm{L}^{\prime}+\bm{L}_{Q}^{\prime}$. The subscripts $\pm$ in Eq. (50) stand for two different “topologies” of the decay. The interference between the (+)(+)( + ) and ()(-)( - ) topologies ensures conservation of $\mathcal{C}$-parity. The angular-momentum coefficients can be written explicitly in terms of Wigner 6-$j$ and 9-$j$ symbols as [38]
@@ V.2 Selection Rules
The transition amplitudes for the decays of a meson into pairs of mesons are obtained by inserting Eq. (51) into Eq. (50). Zeroes of the transition amplitude correspond to selection rules. Zeroes of the angular-momentum coefficients $\xi$ correspond to selection rules associated with conservation of angular momentum. Zeroes of the spatial matrix elements $A$ correspond to selection rules associated with conservation of parity. These selection rules are described in Ref. [23].
For decays into meson pairs with identical spatial wavefunctions, there is an additional selection rule that corresponds to a cancellation in the sum of the two topologies (+)(+)( + ) and ()(-)( - ) in Eq. (50). It was derived in its most general form in Ref. [34] using the symmetries of the matrix elements $A$ for nonrelativistic quark models. The rule states that a quark-antiquark meson is allowed to decay into pairs of mesons with identical spatial wavefunctions only if the $\mathcal{C}\mathcal{P}$-parity of the decaying meson and its total constituent-quark spin $S_{Q}$ stand in the relation $CP=(-1)^{S_{Q}+S_{q}}$, where $S_{q}$ is the total spin of the light-quark pair that is created. Since for nonflip triplet decay models one has $S_{q}=1$, the selection rule becomes $CP=(-1)^{S_{Q}+1}$. According to this selection rule, decays into pairs of mesons with identical spatial wavefunctions are allowed for quarkonia, which have $CP=(-1)^{S_{Q}+1}$, and forbidden for quarkonium hybrids, which have $CP=(-1)^{S_{Q}}$.
Quarkonium hybrids in B-O multiplets associated with bound states in the $\Pi_{u}$ potential, the $\Sigma_{u}^{-}$ potential, or coupled $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials have $CP=(-1)^{S_{Q}}$. The B-O equivalent of two quark-model mesons with identical spatial wavefunctions is two heavy mesons in charge-conjugate HQSS doublets, e.g., two $S$-wave mesons, two $P^{-}$-wave mesons, or two $P^{+}$-wave mesons. Our finding that quarkonium hybrids in B-O multiplets with a $\Sigma_{u}^{-}$ component may have nonzero decay rates into pairs of $S$-wave mesons therefore contradicts the selection rule $CP=(-1)^{S_{Q}+1}$, as first pointed out in Ref. [27]. This observation can be generalized to meson pairs other than $S+S$ by applying the selection rules in Section III.2. Bound states in the $\Sigma_{u}^{-}$ potential or coupled $\Pi_{u}$ and $\Sigma_{u}^{-}$ potentials can have nonzero decay rates into pairs of heavy mesons in charge-conjugate HQSS doublets with $j_{1}=j_{2}\geq\frac{1}{2}$, such as $S+S$, $P^{-}+P^{-}$, or $P^{+}+P^{+}$. Bound states in the $\Pi_{u}$ potential can have nonzero decay rates into pairs of heavy mesons in charge-conjugate HQSS doublets with $j_{1}=j_{2}\geq\frac{3}{2}$, such as $P^{+}+P^{+}$. These results support the assignment $S_{q}=0$ for quark-pair creation models of quarkonium-hybrid decays. In fact, contrarily to the widely used nonflip-triplet decay models with $S_{q}=1$, spin-singlet decay models with $S_{q}=0$ allow decays of quarkonium hybrids into meson pairs with identical spatial wavefunctions [39].